# Coupled Variational Reinforcement Learning for Language Model General Reasoning

### v1

@@ 3.5 Training Model and Data
CoVRL is robust across different base models.
@@
We evaluate CoVRL on Qwen2.5-base and Qwen3-base models ranging from 7B to 14B parameters. Note that for Qwen3-base experiments, we replace <think> </think> with <thinking> </thinking> tags, as Qwen3-base fails to properly utilize <think> </think> tags in prompts while responding well to other similar tags. As shown in Table 2, CoVRL delivers consistent performance improvements across all tested models, demonstrating its robustness across model architectures.
CoVRL learns generalizable reasoning abilities from training data.
@@
We also evaluate CoVRL with different training data compositions. The results in Table 3 reveal that training on both math-only and non-math-only data significantly improves base model performance. Notably, models trained solely on mathematical data demonstrate enhanced performance on non-mathematical reasoning tasks, and similarly, models trained on non-mathematical data improve on mathematical tasks. This indicates that our method enables models to acquire generalizable reasoning capabilities that transfer across different domains.
4 Related Works
Verifier-free Reinforcement Learning
@@ 4 Related Works
While reinforcement learning from verifiable rewards (cui2025processreinforcementimplicitrewards; yu2025dapoopensourcellmreinforcement; yang2025qwen3technicalreport; deepseekai2025deepseekr1incentivizingreasoningcapability) has emerged as a common practice, these approaches are restricted to domains where robust verifiers are available. Recent work explores eliminating external verification dependencies. LaTRO (chen2024languagemodelshiddenreasoners) proposes a framework that formulates reasoning as sampling from a latent distribution. JLB (tang2025verifiablerewardsscalingreinforcement) attempts to scale this approach to data that cannot be easily verified. VeriFree (zhou2025reinforcinggeneralreasoningverifiers) shares a similar framework and proposes using probability directly as both a reward and a coefficient for negative log-likelihood loss. RLPR (yu2025rlprextrapolatingrlvrgeneral) introduces a different reward mechanism using the mean of token rewards. These approaches sample from prior distributions, potentially suffering from limited sample efficiency and reasoning-answer misalignment issues. In contrast, RAVR (lin2025ravrreferenceanswerguidedvariationalreasoning) also utilizes posterior distributions. Compared to RAVR, our approach establishes explicit coupling between prior and posterior through a composite distribution and hybrid sampling strategy.
Self-improving language models
@@
Recent work explores various strategies for enhancing language models by leveraging their own outputs. One line focuses on self-generated training signals, including methods that prompt models to rank their responses for preference learning (yuan2025selfrewardinglanguagemodels), apply iterative DPO (chen2025bootstrappinglanguagemodelsdpo) with implicit rewards (rafailov2024directpreferenceoptimizationlanguage), and employ majority voting for test-time improvements (zuo2025ttrltesttimereinforcementlearning). Alternative approaches construct rewards based on reference answers, such as the verifier-free methods discussed above. Among them, zhou2025variationalreasoninglanguagemodels also introduce posterior distributions and update two separate prior and posterior models using the IWAE paradigm (burda2016importanceweightedautoencoders) with an EM algorithm, utilizing external verifiers for variance reduction. In comparison, we adopt online RL training to optimize a single composite distribution with verifier-free rewards.
@@ 5 Conclusion
In this paper, we introduce Coupled Variational Reinforcement Learning, which enables joint optimization of question-only and answer-guided generation while maintaining theoretical grounding through a coupled optimization framework based on composite distributions. Experimental results demonstrate consistent improvements across tasks, validating the effectiveness of our approach.
@@ Appendix A KL Divergence Estimator
In this section, we provide an intuitive understanding of our KL divergence estimators and derive the specific forms used in our hybrid sampling strategy.
Let $r=\frac{q_{\psi}(z|x,y)}{p_{\phi}(z|x)}$ denote the likelihood ratio between posterior and prior distributions. Our composite distribution becomes:
$p^{\prime}(z|x,y)=\frac{1}{2}p_{\phi}(z|x)+\frac{1}{2}q_{\psi}(z|x,y)=p_{\phi}(z|x)\left(\frac{1}{2}+\frac{1}{2}r\right)$
We need to estimate $D_{\text{KL}}(p^{\prime}(z|x,y)\|p_{\phi}(z|x))$ using importance sampling from either the prior or posterior distribution.
@@
The importance sampling ratio is $\frac{p^{\prime}(z|x,y)}{p_{\phi}(z|x)}=\frac{1}{2}+\frac{1}{2}r$, and our KL estimator becomes:
$D_{\text{KL}}^{\text{prior}}=\left(0.5+0.5r\right)\ln\left(0.5+0.5r\right)-\left(0.5r-0.5\right)$
@@
The importance sampling ratio is $\frac{p^{\prime}(z|x,y)}{q_{\psi}(z|x,y)}=\frac{0.5+0.5r}{r}=\frac{0.5}{r}+0.5$, and our KL estimator becomes:
$D_{\text{KL}}^{\text{posterior}}=\left(\frac{0.5}{r}+0.5\right)\ln\left(0.5+0.5r\right)+\left(\frac{0.5}{r}-0.5\right)$
@@
Appendix B Comparison between Existing Approaches
@@
All methods aim to improve reasoning without external verifiers but differ in their approach. JLB (tang2025verifiablerewardsscalingreinforcement) uses log-probability as reward with fixed answer term weighting. LaTRO (chen2024languagemodelshiddenreasoners) incorporates KL regularization between policy and reference models. VeriFree (zhou2025reinforcinggeneralreasoningverifiers) uses the probability instead of log-probability as reward and also uses it as the weight term for NLL loss. RLPR (yu2025rlprextrapolatingrlvrgeneral) optimizes joint answer-reasoning probability using token-level probabilities as rewards. RAVR (lin2025ravrreferenceanswerguidedvariationalreasoning) samples from posterior distributions and optimizes the corresponding variational lower bound. The core differences among these methods lie in sampling distribution, reward design, and the weighting coefficients for the NLL loss terms.

### v2

@@ 3.1 Setup
3.2 Main Results
Table 1 presents our main experimental results across all benchmarks. Our approach achieves consistent improvements over the base model while outperforming the strongest baseline by 2.3% on average, demonstrating the effectiveness of our coupled optimization framework.
3.3 Training Dynamics
We present the evolution of key metrics over training in Figure 4 to better understand the algorithm. This leads to several observations:
The posterior distribution is effective in providing guidance. Figure 4(a) shows steady reward score improvements throughout training, with the posterior distribution consistently outperforming the prior distribution. This persistent score gap validates our answer-guided sampling strategy and confirms that posterior sampling enables more efficient exploration of high-quality reasoning paths.
CoVRL enhances reasoning capabilities with prolonged chain-of-thought traces. Figure 4(b) demonstrates a stable increase in response length, indicating that the model progressively generates more detailed reasoning processes. This trend suggests that CoVRL successfully encourages elaborate chain-of-thought reasoning with more thorough step-by-step explanations.
Regularization provides stable optimization dynamics. Figures 4(c) and 4(d) show stable downward trends in NLL and KL losses. The decreasing losses indicate improved answer prediction and successful regularization, confirming that our variational objective effectively balances the reconstruction and regularization terms.
3.4 Hybrid Sampling Strategy
We examine the impact of our hybrid sampling strategy by varying the mixing ratio $\alpha$ between prior $p_{\phi}(z|x)$ and posterior $q_{\psi}(z|x,y)$ distributions. Here $\alpha$ represents the probability of sampling from the prior distribution. By default, we set $\alpha=0.5$. Figure 5 demonstrates that low prior sampling probability ($\alpha=0.1$) outperforms high prior sampling probability ($\alpha=0.9$). This highlights the important role of the posterior distribution in our algorithm. When $\alpha=0.9$, the model primarily samples from the prior distribution, and we observe reduced reasoning chain length. We attribute this to the difficulty in improving rewards, leading the model to prioritize minimizing KL loss and generating shorter sequences. When $\alpha=0.1$, reasoning chain length increases, and posterior-dominated sampling achieves better performance than prior-dominated sampling. However, due to training-inference mismatch, performance remains inferior to balanced sampling. Training dynamics under these settings are provided in Appendix D.
@@ 3.5 Training Model and Data
CoVRL is robust across different base models. We evaluate CoVRL on the Qwen2.5-base and Qwen3-base models, ranging from 7B to 14B parameters. Note that for Qwen3-base experiments, we replace <think> </think> with <thinking> </thinking> tags, as Qwen3-base fails to properly utilize <think> </think> tags in prompts while responding well to other similar tags. As shown in Table 2, CoVRL delivers consistent performance improvements across all tested models, with gains ranging from 8.6% to 14.0% across different model sizes, demonstrating its robustness and effective scaling with model capacity.
@@
CoVRL learns generalizable reasoning abilities from training data. We also evaluate CoVRL with different training data compositions. The results in Table 3 reveal that training on both math-only and non-math-only data significantly improves base model performance. Notably, models trained solely on mathematical data demonstrate enhanced performance on non-mathematical reasoning tasks, and similarly, models trained on non-mathematical data improve on mathematical tasks. This indicates that our method enables models to acquire generalizable reasoning capabilities that transfer across different domains, suggesting that CoVRL develops fundamental reasoning patterns rather than domain-specific solutions.
4 Related Work
@@
Verifier-free Reinforcement Learning. While RLVR (Cui et al., 2025; Yu et al., 2025a; Yang et al., 2025; DeepSeek-AI et al., 2025) has emerged as common practice, these approaches require robust verifiers, limiting their applicability to broader reasoning tasks. Recent work such as LaTRO (Chen et al., 2024) eliminates this dependency by formulating reasoning as sampling from latent distributions. Subsequent methods such as JLB (Tang et al., 2025), VeriFree (Zhou et al., 2025a), and RLPR (Yu et al., 2025b) extend this framework using probability-based rewards, but these prior-based sampling methods suffer from limited sample efficiency and reasoning-answer misalignment. RAVR (Lin et al., 2025) instead utilizes posterior distributions for sampling. Our approach establishes explicit coupling between prior and posterior through a composite distribution and hybrid sampling strategy.
@@
Self-improving Language Models. Recent work enhances language models by leveraging self-generated training signals, including preference learning (Yuan et al., 2025), iterative DPO (Chen et al., 2025; Rafailov et al., 2024), and test-time voting (Zuo et al., 2025). Alternative approaches construct rewards from reference answers, such as the verifier-free methods discussed above and Zhou et al. (2025b), which introduce posterior distributions and update separate prior and posterior models using IWAE (Burda et al., 2016) with an EM algorithm. In comparison, we optimize a single composite distribution with verifier-free rewards through online RL training.
@@ 5 Conclusion
In this paper, we introduce Coupled Variational Reinforcement Learning for enhancing language model reasoning without external verifiers. Our method enables joint optimization of question-only and answer-guided generation through composite distributions. By establishing coupling via hybrid sampling, CoVRL balances efficient exploration with inference-time transferability.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Our method improves language model reasoning capabilities without requiring external verifiers, potentially democratizing access to advanced AI reasoning. While this enables beneficial applications in education and scientific discovery, we acknowledge that enhanced reasoning abilities could be misused for generating sophisticated misinformation. We encourage the community to consider safety mechanisms and ethical guidelines when deploying such capabilities.
@@ Appendix A KL Divergence Estimator
In this section, we provide the derivation and intuition for our KL divergence estimators used in CoVRL.
A.1 Derivation via Control Variates and Bregman Divergence
Our KL estimators are derived following the control variate approach proposed by Schulman (2020). The key idea is to construct unbiased, low-variance estimators of KL divergence when we can only sample from one of the distributions.
Reverse KL $D_{\text{KL}}(q\|p)$
When sampling from $q$ to estimate $D_{\text{KL}}(q\|p)=\mathbb{E}_{x\sim q}[\log q(x)-\log p(x)]$, the naive unbiased estimator is $\log r$ where $r=\frac{p(x)}{q(x)}$, but this has high variance as it can be negative even though KL is always non-negative. To reduce variance while maintaining unbiasedness, we employ a control variate: we add a zero-mean term that is negatively correlated with the naive estimator. Since $\mathbb{E}_{x\sim q}[r-1]=0$, we can construct the estimator:
$\log r+\lambda(r-1)$ (17)
which is unbiased for any $\lambda$. Setting $\lambda=1$ yields:
$(r-1)-\log r$ (18)
Forward KL $D_{\text{KL}}(p\|q)$
When sampling from $q$ to estimate $D_{\text{KL}}(p\|q)=\mathbb{E}_{x\sim p}[\log p(x)-\log q(x)]$, we need importance weighting. The naive importance-weighted estimator is $r\log r$ where $r=\frac{p(x)}{q(x)}$. Applying the control variate r1r-1 (which has zero mean under $q$), we obtain:
$r\log r-(r-1)$ (19)
Both forms are always non-negative because they measure the Bregman divergence between $\log(x)$ and its tangent line at $x=1$. Since $\log$ is concave, we have $\log(x)\leq x-1$, ensuring non-negativity. This generalizes to f-divergences via the formula $f(r)-f^{\prime}(1)(r-1)$, which measures the distance between a convex function $f$ and its tangent at $r=1$.
A.2 KL Estimators under Hybrid Sampling
Recall that our composite distribution is defined through token-level mixing:
$p^{\prime}(z_{t}|z_{<t},x,y)=\frac{1}{2}p_{\phi}(z_{t}|z_{<t},x)+\frac{1}{2}q_{\psi}(z_{t}|z_{<t},x,y)$ (20)
Let $r_{t}=\frac{q_{\psi}(z_{t}|z_{<t},x,y)}{p_{\phi}(z_{t}|z_{<t},x)}$ denote the token-level likelihood ratio between posterior and prior distributions. Then:
$p^{\prime}(z_{t}|z_{<t},x,y)=p_{\phi}(z_{t}|z_{<t},x)\left(\frac{1}{2}+\frac{1}{2}r_{t}\right)$ (21)
We need to estimate $D_{\text{KL}}(p^{\prime}(z_{t}|z_{<t},x,y)\|p_{\phi}(z_{t}|z_{<t},x))$ at each token position. In our hybrid sampling strategy, we sample complete sequences from either the prior $p_{\phi}(z|x)$ or posterior $q_{\psi}(z|x,y)$, rather than sampling from the composite distribution $p^{\prime}(z|x,y)$ directly. To obtain unbiased estimates of the KL divergence under this sampling scheme, we employ importance weighting that accounts for the distribution mismatch.
@@ A.2 KL Estimators under Hybrid Sampling
The token-level importance sampling ratio is:
$w_{t}=\frac{p^{\prime}(z_{t}|z_{<t},x,y)}{p_{\phi}(z_{t}|z_{<t},x)}=\frac{1}{2}+\frac{1}{2}r_{t}$ (22)
Since we are estimating $D_{\text{KL}}(p^{\prime}\|p_{\phi})$ (i.e., forward KL), we apply the Bregman divergence formula for this direction: $w\log w-(w-1)$. This gives:
$D^{\text{prior}}_{\text{KL}}$ $=w_{t}\log w_{t}-(w_{t}-1)$ (23)
$=\left(\frac{1}{2}+\frac{1}{2}r_{t}\right)\log\left(\frac{1}{2}+\frac{1}{2}r_{t}\right)-\left(\frac{1}{2}r_{t}-\frac{1}{2}\right)$
@@
The token-level importance sampling ratio is:
$w_{t}=\frac{p^{\prime}(z_{t}|z_{<t},x,y)}{q_{\psi}(z_{t}|z_{<t},x,y)}=\frac{1}{2r_{t}}+\frac{1}{2}$ (24)
To estimate $D_{\text{KL}}(p^{\prime}\|p_{\phi})$, we need $\log p^{\prime}-\log p_{\phi}$. Using importance sampling from $q_{\psi}$:
$D_{\text{KL}}^{\text{posterior}}$ $=w_{t}\log\frac{p^{\prime}(z_{t}|z_{<t},x,y)}{p_{\phi}(z_{t}|z_{<t},x)}+(w_{t}-1)$ (25)
$=\left(\frac{1}{2r_{t}}+\frac{1}{2}\right)\log\left(\frac{1}{2}+\frac{1}{2}r_{t}\right)+\left(\frac{1}{2r_{t}}-\frac{1}{2}\right)$
Note the sign difference in the correction term: $(w_{t}-1)$ for prior sampling versus $(w_{t}-1)$ for posterior sampling. Both forms ensure non-negativity through the Bregman divergence property.
@@
Appendix B Comparison among Existing Approaches
@@
All methods aim to improve reasoning without external verifiers but differ in their approach. JLB (Tang et al., 2025) uses log-probability as reward with fixed answer term weighting. LaTRO (Chen et al., 2024) incorporates KL regularization between policy and reference models. VeriFree (Zhou et al., 2025a) uses the probability instead of log-probability as reward and also uses it as the weight term for NLL loss. RLPR (Yu et al., 2025b) optimizes joint answer-reasoning probability using token-level probabilities as rewards. RAVR (Lin et al., 2025) samples from posterior distributions and optimizes the corresponding variational lower bound. The core differences among these methods lie in sampling distribution, reward design, and the weighting coefficients for the NLL loss terms.